Calculate the potential energy of a 5kg mass when it is (i)3m,(ii)6 above the ground.( g=10N\kg)
1
Expert's answer
2010-11-22T14:34:06-0500
In this case the potential energy can be written as
P = mgh, where m is the mass of an object, g is the qravitational acceleration, h is the height above the ground
i. P = 5*10*5 = 150 J
ii. P = 5*10*6 = 300 J